OpenStudy (anonymous):

Please help with this vectors question Write a vector equation and a scalar equation of the plane. Parallel to the yz-plane and including the point (-1, -2, 5). Answer from the book: [x, y, z] = [-1, -2, 5] + s[0, 0, 1] + t[0, 1, 0] x + 1 = 0 Please show all work! Thanks :)

OpenStudy (kinggeorge):

What's a vector that's perpendicular to the yz plane?

OpenStudy (anonymous):

This is what the textbook question says word for word

OpenStudy (anonymous):

This is what I got so far, i=[1,0,0] Ax+By+Cz+D=0 x+(0)y+(0)z+D=0 x+D=0 (-1)+D=0 D=1 scalar equation: x+1=0 Is this right?

OpenStudy (kinggeorge):

That looks right to me. Now you just need to get the vector equation.

OpenStudy (anonymous):

Yea, can you help me with that?

OpenStudy (kinggeorge):

To do the vector equation, you need to find two vectors in that plane. This is surprisingly easy.

OpenStudy (anonymous):

Please teach me because Im extremely confused...My friend showed me how to do it bu I dont understand, here Ill post it,

OpenStudy (anonymous):

Its part b)

OpenStudy (kinggeorge):

The vector that's perpendicular to the yz plane is just <1,0,0>, and there are 3 commonly used vectors that are all perpendicular to each other. These vectors are <1,0,0>, <0,1,0>, and <0,0,1>. If you're given one of these vectors, and asked for another perpendicular vector, choose another one of these.

OpenStudy (kinggeorge):

Thus, you can use <0,1,0> and <0,0,1> for the two vectors needed in your vector equation. As for the starting point, use (-1,2,5) since the plane must contain that line.

OpenStudy (anonymous):

So, if I copy that sheet, is it correct?

OpenStudy (anonymous):

Oh and by the way its [-1, -2, 5]

OpenStudy (kinggeorge):

Almost. That sheet gives you one possible vector. You need to choose another value for y and z, and find a vector from (-1,-2,5) to that second point. Multiply this second vector by s, and add it to the equation.

OpenStudy (anonymous):

Can you help me with that?

OpenStudy (kinggeorge):

Pick another coordinate with x=1.

OpenStudy (anonymous):

Im sorry lol but I dont know what you mean

OpenStudy (kinggeorge):

(1, y, z) choose some values for y and z.

OpenStudy (anonymous):

Any values?

OpenStudy (kinggeorge):

any. although small integers are usually nice.

OpenStudy (anonymous):

ok, how about [1, 3, 4]

OpenStudy (kinggeorge):

Did I say x=1? I keep forgetting negative signs. It should be -1. So we'll go with (-1, 3, 4).

OpenStudy (anonymous):

Ok

OpenStudy (kinggeorge):

Now, what's the vector from (-1, -2, 5) to (-1, 3, 4)?

OpenStudy (anonymous):

umm is it, [-1 +1, 3+2, 4-5] [0, 5, -1] ?

OpenStudy (kinggeorge):

Bingo. So now you have the second vector you need.

OpenStudy (kinggeorge):

Using the stuff written on the sheet, your vector equation would then be <-1,-2,5>+t<0,2,-5>+s<0,5,-1>

OpenStudy (anonymous):

Are you sure thats how it is?

OpenStudy (kinggeorge):

Just like the other problem I did earlier, there are multiple vector equations for this one plane.

OpenStudy (kinggeorge):

The "simplest" one is just the solution that the book gave you.

OpenStudy (anonymous):

Ok, THANKS A TON!! :) <3 I have one more question similar to this if you'd like to help :P

OpenStudy (anonymous):

@KingGeorge Is it like this? [x, y, z]=[-1, -2, 5] + s[0, 5, -1] + t[0, 2, -5]

OpenStudy (kinggeorge):

That is also correct.

OpenStudy (anonymous):

Ok thanks!